OpenStudy (anonymous):

What is the limit of (1-cos(h))^2 / h as h approaches zero?

OpenStudy (anonymous):

Do you know the limit \[\lim_{h\to0}\frac{1-\cos h}{h}~~~?\] You can rewrite your limit as \[\lim_{h\to0}\frac{(1-\cos h)^2}{h}=\lim_{h\to0}\frac{1-\cos h}{h}\cdot\lim_{h\to0}~(1-\cos h)\]

OpenStudy (anonymous):

If you're not familiar with that first limit, you could evaluate it using the squeeze theorem.

OpenStudy (anonymous):

Thanks. I completely forgot about rewriting the function.

OpenStudy (anonymous):

You're welcome